Jay invested a sum of money at 6% per year and 3 times as much at 4% per year. How much did he invest at the rate if his annual return totaled $144
Answer by shree840(260)

annual return will mean period is one year.
x(1.06)+3x(1.04)=144
or 1.06x+3.12x=144
4.18x+144
x=144/4.18
x=34.44